Grade 128 grade maths

Factorise the following expressions
5x^2y-15xy^2
Factorise the following expressions
5x^2y-15xy^2

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

5x^2y = 5×x × x×y
15xy^2= 3×5×x × y ×y
The commom factors are 5, x and y
5x^2y – 15xy^2= (5×x × x×y)–(3×5×x × y ×y)
=5×x × y[ x – (3 × y ) ]
= 5xy ( x-3y)

Thanks